OpenStudy (anonymous):

7 0 How would you graph y=x^3

OpenStudy (precal):



OpenStudy (anonymous):

why

OpenStudy (anonymous):

how did u do that

OpenStudy (precal):

It is a parent function. You should memorize them if you get a chance. y=x linear y=x^2 quadratic y=x^3 cubic you can also create a table to graph it

OpenStudy (precal):